Three charges Q1, Q2 , and Q3 are located in a strCan anyone help me with the third part?? Thanks so much!
Answer #1

Q1 -Q2 X12 x23

The figure show the arrangement of the charges. As you can see we want to find x. The net force on Q3 is

F=k_e\frac{Q_1Q_3}{x^2}+k_e\frac{-Q_2Q_3}{x_{23}\,^2}

Like we do not know the distance from Q2 to Q3 we go tu use the relation

x=x_{12}+x_{23}

x_{23}=x-x_{12}

Then

F=k_e\frac{Q_1Q_3}{x^2}+k_e\frac{-Q_2Q_3}{(x-x_{12})\,^2}

F=\left(\frac{Q_1}{x^2}-\frac{Q_2}{(x-x_{12})\,^2} \right )k_eQ_3

Like the net force must be zero

0=\left(\frac{Q_1}{x^2}-\frac{Q_2}{(x-x_{12})\,^2} \right )k_eQ_3

0=\frac{Q_1}{x^2}-\frac{Q_2}{(x-x_{12})\,^2}

0=\frac{Q_1(x-x_{12})\,^2-Q_2(x^2)}{x^2(x-x_{12})\,^2}

0={Q_1(x-x_{12})\,^2-Q_2(x^2)}

0=Q_1(x^2-2x_{12}x+x_{12}\,^2)-Q_2x^2

0=Q_1x^2-2Q_1x_{12}x+Q_1x_{12}\,^2-Q_2x^2

0=(Q_1-Q_2)x^2-2Q_1x_{12}x+Q_1x_{12}\,^2

The solution of this equation is given by

x=\frac{-(-2Q_1x_{12})\pm\sqrt{(-2Q_1x_{12})^2-4(Q_1-Q_2)(Q_1x_{12}\,^2)}}{2(Q_1-Q_2)}

x=\frac{2Q_1x_{12}\pm\sqrt{4Q_1\,^2x_{12}\,^2-4Q_1\,^2x_{12}\,^2+4Q_2Q_1x_{12}\,^2}}{2(Q_1-Q_2)}

x=\frac{2Q_1x_{12}\pm\sqrt{4Q_2Q_1x_{12}\,^2}}{2(Q_1-Q_2)}

x=\frac{2Q_1x_{12}\pm2\sqrt{Q_2Q_1}x_{12}}{2(Q_1-Q_2)}

x=\frac{Q_1\pm\sqrt{Q_2Q_1}}{Q_1-Q_2}x_{12}

x=\frac{(1.56\times 10^{-6}C)\pm\sqrt{(2.84\times 10^{-6}C)(1.56\times 10^{-6}C)}}{(1.56\times 10^{-6}C)-(2.84\times 10^{-6}C)}(0.235m)

Here we have two solutions

x_1=-0.673m  at the left of the charge 1

x_2=0.100m at the right of the charge 1